### v1

# MixNet: Towards Effective and Efficient UHD Low-Light Image Enhancement

Abstract
With the continuous advancement of imaging devices, the prevalence of Ultra-High-Definition (UHD) images is rising. Although many image restoration methods have achieved promising results, they are not directly applicable to UHD images on devices with limited computational resources due to the inherently high computational complexity of UHD images. In this paper, we focus on the task of low-light image enhancement (LLIE) and propose a novel LLIE method called MixNet, which is designed explicitly for UHD images. To capture the long-range dependency of features without introducing excessive computational complexity, we present the Global Feature Modulation Layer (GFML). GFML associates features from different views by permuting the feature maps, enabling efficient modeling of long-range dependency. In addition, we also design the Local Feature Modulation Layer (LFML) and Feed-forward Layer (FFL) to capture local features and transform features into a compact representation. This way, our MixNet achieves effective LLIE with few model parameters and low computational complexity. We conducted extensive experiments on both synthetic and real-world datasets, and the comprehensive results demonstrate that our proposed method surpasses the performance of current state-of-the-art methods. The code will be available at https://github.com/zzr-idam/MixNet.
1 Introduction
In recent years, high-resolution imaging has experienced significant advancements due to the emergence of sophisticated imaging sensors and displays. This has led to a rapid evolution of Ultra-High-Definition (UHD) imaging technology. However, the inherent high-resolution characteristics of UHD images make them more susceptible to noise during the imaging process. Additionally, the increased number of pixels in UHD images challenges the efficiency of existing image processing methods. These undoubtedly affect the quality of UHD imaging.
In this paper, we focus on the task of UHD low-light image enhancement (LLIE). Images captured under low-light conditions usually exhibit significant degradation, including reduced visibility, low contrast, and high noise levels. Recovering ideal normal-light images from UHD low-light images is a long-studied problem.
Figure 1: Model complexity and performance comparison between our proposed MixNet and other LLIE methods on the UHD-LOL4K dataset. Since the inability of most methods to process UHD images directly, the calculation of FLOPs is performed based on the scale of 1024 $\times$ 1024 images (the maximum resolution that these algorithms can process on a single GPU). Our MixNet achieves a better trade-off between model complexity and performance.
Although many learning-based LLIE methods have made significant progress, they always focus on low-resolution images Wei et al. (2018); Zhao et al. (2021); Wu et al. (2022); Xu et al. (2022); Cai et al. (2023). For example, in the LLIE task, the most common dataset used is LOL Wei et al. (2018), which consists of images with resolutions below 1K ($1920\times 1080$). To achieve better results, current state-of-the-art (SOTA) methods Hou et al. (2023); Cai et al. (2023) commonly employ larger models and introduce a significant number of parameters. When faced with UHD images, these SOTA methods often encounter memory overflow and cannot perform full-resolution inference on consumer-grade GPUs. With the prevalence of UHD images, designing a LLIE algorithm tailored explicitly for UHD images becomes increasingly essential.
[…]
In this paper, we propose a simple but effective UHD-LLIE method, called MixNet. Specifically, we design the Feature Mixing Block which mainly contains Global Feature Modulation Layer (GFML), Local Feature Modulation Layer (LFML), and Feed-forward Layer (FFL). GFML is the key component of MixNet, enabling efficient modeling of long-range dependency. The use of self-attention in transformers endows them with powerful long-range dependency modeling capabilities Vaswani et al. (2017). However, this approach comes with significant computational overhead. Inspired by the vanilla MLP-Mixer Tolstikhin et al. (2021), but different from it, GFML encodes the feature maps of images from the perspectives of the width $W$, height $H$, and channel $C$. Through simple dimensional transformation operations, the information encoded in each feature map is associated and fused. It is worth noting that dimension transformation operations do not introduce additional parameters, which means that by permuting the feature maps from different views, GFML achieves efficient modeling of long-range dependency. LFML is designed to capture the local features of images. LFML adaptively calculates a set of parameters for the redistribution of weights in the channel dimension, emphasizing the importance of specific channels. The results of GFML and LFML are fed to FFL which aim to transform features to a compact representation. This way, our MixNet achieves a better trade-off between model complexity and performance. Extensive experiments on both synthetic and real-world datasets demonstrate that MixNet surpasses existing SOTA methods in terms of restoration quality and generalization ability.
Our contributions can be summarized below:
  • We propose a simple but effective UHD-LLIE method called MixNet. MixNet can perform full-resolution inference of UHD images on consumer-grade GPUs. Compared to patch-based strategies, full-resolution inference can avoid the occurrence of boundary artifacts.
  • We develop a novel efficient feature modulation mechanism to capture long-range dependency, which does not need to project the input image from 3D to 2D; it maintains the spatial properties between pixels by modeling directly on 3D space.
  • We evaluate the proposed method quantitatively and qualitatively on synthetic and real-world datasets, and the results show that our MixNet achieves a favorable trade-off between accuracy and model complexity. Furthermore, MixNet shows excellent scalability and flexibility, performing well on other image restoration tasks (e.g., dehazing).
2 Related Work
2.1 Low-Light Image Enhancement
Learning-based methods have been successfully applied to the LLIE task Fu et al. (2022); Guo et al. (2020); Zhou et al. (2022); Wu et al. (2022); Xu et al. (2022). RetinexNet Wei et al. (2018) relies on the Retinex theory and designs a Decom-Net for decomposition and an Enhance-Net for illumination adjustment. Transformers have gained significant traction in computer vision since ViT Dosovitskiy et al. (2020) employs transformers for visual recognition tasks. Restormer Zamir et al. (2022) efficiently achieves image restoration by calculating self-attention at the channel dimension, thereby avoiding significant computational overhead. It demonstrates excellent performance across multiple tasks. Retinexformer Cai et al. (2023) introduces Retinex theory Land (1977) to the transformer and designs an illumination-guided transformer that utilizes illumination representations to direct the modeling of non-local interactions of regions with different lighting conditions. Diffusion models show realistic and detailed image generation through a sequence of denoising refinements. PyDiff Zhou et al. (2023) introduces diffusion models to LLIE to recover realistic details and performs sampling in a pyramid resolution style to get a fast generation. Although the methods above have achieved promising results in the LLIE task, they are not directly applicable to UHD images because the additional pixels of UHD images result in high computational overhead.
Figure 2: The overall architecture of the proposed MixiNet. MixNet first transforms the input low-light image into the feature space using a DownSampler, performs feature extraction using a series of Feature Mixing Blocks (FMBs), and then reconstructs these extracted features using an UpSampler. The FMB mainly contains a Global Feature Modulation Layer (GFML), a Local Feature Modulation Layer (LFML), and a Feed-forward Layer (FFL).
2.2 UHD Image Restoration
UHD image restoration has been an emerging topic in recent years Liang et al. (2021); Zamfir et al. (2023); Guan et al. (2022). In  Zheng et al. (2021a, b), the authors introduced bilateral learning to achieve dehazing of the UHD image and HDR reconstruction of the UHD image. The core idea is to learn the local affine coefficients of the bilateral grid from the low-resolution image and apply them to the full-resolution image. UHD-SFNet Wei et al. (2022) and FourUHD Li et al. (2023) explore underwater UHD image enhancement and UHD low-light image enhancement in the Fourier domain. They are all inspired by the discovery of most luminance information concentrates on amplitudes. Due to the unavoidable need for downsampling in existing UHD image restoration methods, NSEN Yu et al. (2023) proposes a spatial-variant and invertible nonuniform downsampler that adaptively adjusts the sampling rate according to the richness of details. LLFormer Wang et al. (2023) is the first transformer-based method to complete UHD-LLIE task, but like most previous methods, it is unable to perform full-resolution inference on consumer-grade GPUs. Despite the promising results achieved by the aforementioned approaches, there is still room for a favorable trade-off between reconstruction performance and model efficiency.
3 Methodology
MixNet aims to recover ideal UHD normal-light images from low-light images which often exhibit significantly reduced visibility, low contrast, and high levels of noise. We provide the overall pipeline of our method and further details on the critical components of our approach below.
3.1 Overview
An overview of MixNet is shown in Fig. 2. For a given low-light image ${x}\in\mathbb{R}^{H\times W\times 3}$, we first map it to the feature space through a DownSampler to obtain the low-level features ${F}_{0}\in\mathbb{R}^{H\times W\times C}$, where $H$, $W$, and $C$ represent height, width, and channel, respectively. Then, the multiple stacked Feature Mixing Blocks (FMBs) are used to generate finer deep features ${F}_{i}$ from ${F}_{0}$ for normal-light image reconstruction, where a FMB has a Global Feature Modulation Layer (GFML), a Local Feature Modulation Layer (LFML) and a Feed-forward Layer (FFL). Lastly, the sum of the final features ${F}_{f}$ and ${F}_{0}$ is fed to a UpSampler to obtain the restored image ${\hat{y}}$.
Figure 3: The overall architecture of the proposed (a) Global Feature Modulation Layer, (b) Local Feature Modulation Layer, and (c) Feed-forward Layer.
Figure 4: Visualization of dimensional transformation operations. MixNet can capture long-range dependency of features with few parameters by employing simple dimension transformation operations.
3.2 Global Feature Modulation Layer
Recent studies suggest that the notable performance of transformers across diverse tasks stems from their implementation of the key multi-head self-attention (MHSA) mechanism Vaswani et al. (2017). MHSA enables the model with the capability of long-range feature interaction and dynamic spatial weighting, both of which contribute to get promising results. However, this mechanism comes with significant computational overhead. To capture the long-range dependency of features, we ingeniously establish long-range relationships in both spatial and channel dimensions by employing simple dimension transformation operations. This allows for achieving long-range dependency of features by only using a few parameters.
The architecture of the GFML is shown in Fig. 3 (a) and the more details of dimension transformation operations are shown in Fig. 4. Specifically, we first adjust the resolution of the normalized input features, and then perform some dimension transformation operations on them. Given the input feature $F_{in}$, this procedure can be formulated as:
$F_{t}$ $=\operatorname{IP}(F_{in}),$ (1)
$\hat{F}_{t}$ $=\underbrace{\operatorname{Permute}(\operatorname{GELU}(\operatorname{Conv_{1\times 1}}(F_{t})))}_{\times 3},$ (2)
where $F_{t}$ and $\hat{F}_{t}$ are intermediate results. $\operatorname{IP}(\cdot)$ corresponds to the interpolation operation, $\operatorname{Conv_{1\times 1}}(\cdot)$ is a $1\times 1$ convolution, $\operatorname{GELU}(\cdot)$ represents GELU function, $\operatorname{Permute}(\cdot)$ denotes the dimension transformation operation and ×$\times 3$ denotes three operations in sequence. It is worth noting that before executing the last dimension transformation operation, we replace the GELU function with the Sigmoid function.
Afterward, We use interpolation to adjust the feature $\hat{F}_{t}$ to its original resolution to estimate the attention map and adaptively modulate the input $F_{in}$ according to the estimated attention via element-wise product. This process can be written as:
$F_{g}=\operatorname{IP}(\hat{F}_{t})\odot F_{in},$ (3)
where $F_{g}$ are the final output features and $\odot$ represents element-wise product.
3.3 Local Feature Modulation Layer
LFML aims to capture the local feature of images and the architecture of it is shown in Fig. 3 (b). For a set of feature maps, their impact on the final results is different Zhang et al. (2018b), and we hope the model can focus on the important feature maps. We first feed the normalized input features into global average pooling to shift the focus of the model from space dimension to channel dimension, and then feed them into a series of convolutions to get deep features. Finally, these deep features are processed by the sigmoid function to obtain the weight of each channel. The normalized input features adaptively adjust the importance of the channel based on these weights via element-wise product. The LFML can be written as:
$F_{w}$ $=\operatorname{Sigmoid}(\operatorname{Conv}(\operatorname{ReLU}(\operatorname{Conv}(\operatorname{GAP}(F_{in}))))),$ (4)
$F_{l}$ $=F_{w}\odot F_{in},$ (5)
where $F_{in}$ stand for the input features, $F_{w}$ are the channel weights and $F_{l}$ denote the final output features. $\operatorname{Sigmoid}(\cdot)$ represents Sigmoid function, $\operatorname{Conv}(\cdot)$ is a $1\times 1$ convolution and $\operatorname{GAP}(\cdot)$ denotes global average pooling.
3.4 Feed-forward Layer
To transform features into a compact representation, we introduce FFL into our model. As shown in Fig. 3 (c), FFL comprises a $3\times 3$ convolution, a $1\times 1$ convolution, and a GELU function. Within this, the first $3\times 3$ convolution encodes the spatially local contexts and doubles the number of channels of the input features for mixing channels; the later $1\times 1$ convolution reduces the channels back to the original input dimension. In this way, the model learns a more compact representation of features and ignores some unimportant information Tan and Le (2021). The FFL can be expressed as:
$F_{out}=\operatorname{Conv_{1\times 1}(\operatorname{GELU}(\operatorname{Conv}_{3\times 3}((F_{in})))))},$ (6)
where $F_{in}$ and $F_{out}$ are the input features and output features, respectively.
3.5 Feature Mixing Block
Here, we describe our FMB pipeline in general, which can be written as:
$\hat{F}_{in}$ $=\operatorname{Conv}[\operatorname{GFML}(\operatorname{LN}(F_{in}));\operatorname{LFML}(\operatorname{LN}(F_{in}))]+F_{in},$ (7)
$F_{out}$ $=\operatorname{FFL}(\operatorname{LN}(\hat{F}_{in}))+\hat{F}_{in},$ (8)
where $F_{in}$ stand for the input features, $\hat{F}_{in}$ are the intermediate features and $F_{out}$ denote the output features. [;] is concatenation operation, and $\operatorname{Conv}$ respresents the $1\times 1$ convolution.
3.6 Loss Functions
To optimize the weights and biases of the network, we utilize the L1 loss in the RGB color space as the basic reconstruction loss:
$L=\|y-\hat{y}\|_{1},$ (9)
where $y$ and $\hat{y}$ denote the ground truth and restored image, respectively.
Figure 5: Visual quality comparisons. The image can be zoomed in for better visualization. The last row shows the color histogram of the image. Our method has the closest color to GT image.
Table 1: Comparison of quantitative results on UHD-LOL4K and UHD-LL datasets. Best and second best values are indicated with bold text and underlined text respectively.
Type Methods UHD-LOL4K UHD-LL Average Parameter
PSNR SSIM LPIPS PSNR SSIM LPIPS PSNR SSIM LPIPS
non-UHD Zero-DCE Guo et al. (2020) 17.19 0.850 0.193 17.08 0.663 0.513 17.14 0.757 0.353 79.416K
Zero-DCE++ Li et al. (2021) 15.58 0.835 0.222 16.41 0.630 0.530 16.00 0,733 0.376 10.561K
RUAS Liu et al. (2021) 14.68 0.758 0.274 13.56 0.749 0.460 14.12 0.754 0.367 3.438K
Uformer Wang et al. (2022) 29.99 0.980 0.034 19.28 0.849 0.356 24.64 0.915 0.195 20.628M
Restormer Zamir et al. (2022) 36.91 0.988 0.023 22.25 0.871 0.289 29.58 0.930 0.156 26.112M
UHD LLFormer Wang et al. (2023) 37.33 0.989 0.020 22.79 0.853 0.264 30.06 0.921 0.142 13.152M
UHDFour Li et al. (2023) 36.12 0.990 0.021 26.22 0.900 0.239 31.17 0.945 0.130 17.537M
MixNet 39.22 0.992 0.006 27.54 0.862 0.134 33.38 0.927 0.070 7.770M
4 Experiment
4.1 Implementation Details and Datasets
4.1.1 Implementation Details
We conduct experiments in PyTorch on six NVIDIA GeForce RTX 3090 GPUs. To optimize the network, we employ the Adam optimizer with a learning rate $2\times 10^{-4}$. We randomly crop the full-resolution 4K image to a resolution of $512\times 512$ as the input and perform 300k iterations of training with a batch size of 24. To augment the training data, random horizontal and vertical flips are applied to the input images. The number of FMB and feature channels is set to 8 and 48, respectively. DownSampler and UpSampler are both composed of a sub-pixel convolutional layer Shi et al. (2016) and a $3\times 3$ convolutional layer.
4.1.2 Datasets
UHD-LOL.
UHD-LOL Wang et al. (2023) is a large-scale benchmark dataset containing UHD images in low-light conditions and includes two subsets, UHD-LOL4K and UHD-LOL8K, containing 4K and 8K-resolution images, respectively. In this paper, we use the UHD-LOL4K to verify the effectiveness of MixNet. UHD-LOL4K contains 8,099 image pairs, 5,999 for training and 2, 100 for testing.
UHD-LL.
UHD-LL Li et al. (2023) is a real-world low-noise/normal-clear paired image dataset that contains 2,150 pairs of 4K UHD data saved in 8bit sRGB format. It has been split into two parts: 2,000 pairs for training and 115 pairs for testing.
LOL.
In addition to the aforementioned UHD datasets, we also evaluate our method on the LOL dataset Wei et al. (2018) which is the commonly used dataset in the LLIE task. LOL has 500 image pairs and the training and testing sets are split in proportion to 485:15.
Figure 6: Visual quality comparisons. The image can be zoomed in for improved visualization. Our method can effectively remove the noise and reconstruct well-exposed image details.
4.2 Comparisons with the State-of-the-art Methods
4.2.1 Quantitative Results
We employ three well-known image quality assessment metrics, PSNR, SSIM Wang et al. (2004) and LPIPS (Alex version) Zhang et al. (2018a) to quantify the performance of different methods. PSNR and SSIM are calculated along the RGB channels. For UHD datasets, we compare our method with Zero-DCE Guo et al. (2020), Zero-DCE++ Li et al. (2021), RUAS Liu et al. (2021), Uformer Wang et al. (2022), Restormer Zamir et al. (2022), LLFormer Wang et al. (2023) and UHDFour Li et al. (2023). It is worth noting that most methods cannot directly perform full-resolution inference on UHD images. Therefore, for these methods, we adopt a strategy of dividing the images into multiple patches without overlapping for interference, and then the results are stitched together. For the LOL dataset, in addition to the mentioned methods, we also compare our MixNet with methods RetinexNet Wei et al. (2018), KinD Zhang et al. (2019), DSLR Lim and Kim (2020), RUAS Liu et al. (2021), ELGAN Jiang et al. (2021) and Retinexformer Cai et al. (2023).
In Tab. 1, we report the quantitative comparison results among different methods on the UHD dataset. When compared with the previous SOTA method FourUHD, our MixNet gains by an average of 2.21dB on the UHD-LOL4K and UHD-LL datasets. Compared to other methods specifically designed for UHD images, our MixNet achieves impressive results with few parameters. Our method also performs well on the LOL dataset and the results are reported in Tab. 2. It is worth noting that the resolution of LOL images is relatively low. Therefore, we replace the sub-pixel convolutional layer in DownSampler and UpSampler with the Residual Block of comparable parameters.
All these results clearly suggest the outstanding effectiveness and efficiency advantage of our MixNet.
4.2.2 Qualitative Results
In addition to the quantitative evaluations, we also provide qualitative comparisons of the proposed method. Figs. 5 and  6 present the visual comparisons. MixNet can effectively enhance the poor visibility and low contrast or low-light regions, reliably remove the noise without introducing spots and artifacts, and robustly preserve the color.
Table 2: Comparison of quantitative results on LOL dataset. Best and second best values are indicated with bold text and underlined text respectively.
Type Methods PSNR SSIM
non-UHD RetinexNet Wei et al. (2018) 16.77 0.425
KinD Zhang et al. (2019) 17.65 0.772
DSLR Lim and Kim (2020) 14.98 0.596
Z_DCE Guo et al. (2020) 14.86 0.562
Z_DCE++ Li et al. (2021) 14.75 0.518
RUAS Liu et al. (2021) 16.40 0.503
ELGAN Jiang et al. (2021) 17.48 0.652
Uformer Wang et al. (2022) 18.55 0.721
Restormer Zamir et al. (2022) 22.37 0.816
Retinexformer Cai et al. (2023) 25.16 0.845
UHD UHDFour Li et al. (2023) 23.09 0.871
LLFormer Wang et al. (2023) 23.65 0.816
MixNet 24.21 0.860
4.3 Ablation Study
We further conduct extensive ablation studies to better understand and evaluate each component in the proposed MixNet. For fair comparisons with the designed baselines, we implement all experiments with the same setting on UHD-LOL4K dataset. We gradually replace the proposed module with the Residual Block (ResBlock) He et al. (2016) of comparable parameters, and the experimental results in Tab. 4.
Based on all the results in the Tab. 4, all the key designs contribute to the best performance of the full model. Without the GFML (#a, #c, #d, #f), the model shows the most severe performance degradation. This aligns with our hypothesis, as the excellent long-range dependency modeling capability brought by GMFL is one of the key factors contributing to the outstanding performance of MixNet. And the result suggests the importance of long-range dependency modeling among features. From the results of #g and #e, we find that the impact of FFL on performance is higher than that of GFML. We believe that this might be the ability of the FFL to learn compact feature representations, which the ResBlock lacks. And any CNN is capable of learning local features. From the results of #c and #d, we observe that when using the block we proposed alone, LFML has a higher impact on the results than FFL. This may be because when the feature representation is not effective enough (not from GFML and LFML), learning its compact representation has little effect. We also provide visualizations of the feature maps corresponding to #e, #f, #g and #h. As shown in Fig. 7, it is evident that our method reconstructs feature maps with sharper textures.
Table 3: Comparison of quantitative results on O-Haze dataset. Best and second best values are indicated with bold text and underlined text respectively.
Task Methods PSNR SSIM
non-UHD DCP He et al. (2010) 14.54 0.333
BCCR Meng et al. (2013) 15.20 0.361
CAP Zhu et al. (2015) 15.35 0.342
NLD Berman et al. (2016) 14.76 0.396
DehazeNet Cai et al. (2016) 12.39 0.359
AodNet Li et al. (2017) 15.87 0.581
GridNet Liu et al. (2019) 17.44 0.603
PMSNet Chen et al. (2019) 16.99 0.597
MSBDN Dong et al. (2020) 17.46 0.635
UHD MGBL Zheng et al. (2021a) 18.72 0.614
NSEN Yu et al. (2023) 19.23 0.672
MixNet 21.46 0.650
4.4 Extension Experiments
Although we primarily focus on the UHD-LLIE task in this paper, it is important to note that we do not introduce specific prior knowledge related to low-light conditions. Therefore, our method can be considered as a general approach that performs well not only in UHD-LLIE task but also in various UHD image restoration task, such dehazing.
4.4.1 Datasets
We conduct dehazing experiments on the O-HAZE dataset which has 45 paired hazy and haze-free images. Following the prior work Yu et al. (2023), we randomly choose 35 images for training and the others for testing.
4.4.2 Results
We compare the proposed MixNet with several image dehazing approaches Berman et al. (2016); He et al. (2010); Meng et al. (2013); Zhu et al. (2015); Li et al. (2017); Dong et al. (2020); Liu et al. (2019); Cai et al. (2016); Chen et al. (2019); Zheng et al. (2021a); Yu et al. (2023). Results are reported in Tab. 3. MixNet achieves a substantial gain of over 2 dB compared to NSEN in PSNR on the UHD image dehazing task.
Table 4: Ablation study of proposed blocks. In order to ensure consistency in parameter magnitude, we gradually replace the proposed module with Residual Block of comparable parameters.
# GFML LFML FFL PSNR SSIM Parameter
(a) 30.73 0.968 7.178M
(b) 34.16 0.985 7.475M
(c) 33.93 0.984 6.882M
(d) 32.52 0.976 7.769M
(e) 36.44 0.990 8.066M
(f) 34.47 0.987 7.474M
(g) 35.80 0.988 7.179M
(h) 39.22 0.992 7.770M
Figure 7: Visualizations of the feature maps before the UpSampler. The image used here is the low-light image shown in Fig. 6. Our method reconstructs feature maps with sharper textures.
5 Conclusion
In this paper, we proposed a simple yet efficient framework called MixNet to solve the efficient UHD-LLIE problem. MixNet has a series of FMBs which mainly contain a GFML, LFML, and FFL. GFML explored long-range dependency modeling upon the simple dimensional transformation operation. By permuting the feature maps, GFML associates features from different views. LFML and FFL are designed to capture local features and transform features into a compact representation, respectively. Extensive experiments on both synthetic and real-world datasets showed that the proposed MixNet is more efficient than state-of-the-art methods while achieving competitive performance.

### v2

@@
-MixNet: Towards Effective and Efficient UHD Low-Light Image Enhancement
+MixNet: Efficient Global Modeling for Ultra-High-Definition Image Restoration
 Abstract
-With the continuous advancement of imaging devices, the prevalence of Ultra-High-Definition (UHD) images is rising.
-Although many image restoration methods have achieved promising results, they are not directly applicable to UHD images on devices with limited computational resources due to the inherently high computational complexity of UHD images.
-In this paper, we focus on the task of low-light image enhancement (LLIE) and propose a novel LLIE method called MixNet, which is designed explicitly for UHD images.
+Recent advancements in image restoration methods employing global modeling have shown promising results.
+However, these approaches often incur substantial memory requirements, particularly when processing ultra-high-definition (UHD) images.
+In this paper, we propose a novel image restoration method called MixNet, which introduces an alternative approach to global modeling approaches and is more effective for UHD image restoration.
 To capture the long-range dependency of features without introducing excessive computational complexity, we present the Global Feature Modulation Layer (GFML).
@@
 In addition, we also design the Local Feature Modulation Layer (LFML) and Feed-forward Layer (FFL) to capture local features and transform features into a compact representation.
-This way, our MixNet achieves effective LLIE with few model parameters and low computational complexity.
-We conducted extensive experiments on both synthetic and real-world datasets, and the comprehensive results demonstrate that our proposed method surpasses the performance of current state-of-the-art methods.
-The code will be available at https://github.com/zzr-idam/MixNet.
-1 Introduction
+This way, our MixNet achieves effective restoration with low inference time overhead and computational complexity.
+We conduct extensive experiments on four UHD image restoration tasks, including low-light image enhancement, underwater image enhancement, image deblurring and image demoiréing, and the comprehensive results demonstrate that our proposed method surpasses the performance of current state-of-the-art methods.
+The code will be available at https://github.com/5chen/MixNet.
+Introduction
 In recent years, high-resolution imaging has experienced significant advancements due to the emergence of sophisticated imaging sensors and displays.
@@
 Additionally, the increased number of pixels in UHD images challenges the efficiency of existing image processing methods.
-These undoubtedly affect the quality of UHD imaging.
-In this paper, we focus on the task of UHD low-light image enhancement (LLIE).
-Images captured under low-light conditions usually exhibit significant degradation, including reduced visibility, low contrast, and high noise levels.
-Recovering ideal normal-light images from UHD low-light images is a long-studied problem.
-Figure 1: Model complexity and performance comparison between our proposed MixNet and other LLIE methods on the UHD-LOL4K dataset.
-Since the inability of most methods to process UHD images directly, the calculation of FLOPs is performed based on the scale of 1024 $\times$ 1024 images (the maximum resolution that these algorithms can process on a single GPU).
-Our MixNet achieves a better trade-off between model complexity and performance.
-Although many learning-based LLIE methods have made significant progress, they always focus on low-resolution images Wei et al.
-(2018); Zhao et al.
-(2021); Wu et al.
-(2022); Xu et al.
-(2022); Cai et al.
-(2023).
-For example, in the LLIE task, the most common dataset used is LOL Wei et al.
-(2018), which consists of images with resolutions below 1K ($1920\times 1080$).
-To achieve better results, current state-of-the-art (SOTA) methods Hou et al.
-(2023); Cai et al.
-(2023) commonly employ larger models and introduce a significant number of parameters.
+Therefore, recovering clear, realistic, and clean images from degraded UHD images (e.g., low-light, blur) is an extremely challenging task.
+Although many learning-based restoration methods have made significant progress, they always focus on low-resolution images (Zamir et al.
+2022; Wang et al.
+2022; Yang et al.
+2023).
+For example, in the LLIE task, the most common dataset used is LOL (Wei et al.
+2018), which consists of images with a resolution of $600\times 400$.
 When faced with UHD images, these SOTA methods often encounter memory overflow and cannot perform full-resolution inference on consumer-grade GPUs.
-With the prevalence of UHD images, designing a LLIE algorithm tailored explicitly for UHD images becomes increasingly essential.
-[…]
-In this paper, we propose a simple but effective UHD-LLIE method, called MixNet.
+This limitation may potentially impact the applications of UHD imaging devices.
+Figure 1: Our MixNet significantly outperforms state-of-the-art UHD image restoration methods, including UHD, UHDFour, UHDformer and ESDNet, on four UHD image restoration benchmarks.
+Recently, there have been some explorations into restoration methods for UHD images (Wang et al.
+2023; Li et al.
+2023a; Wang et al.
+2024; Zheng et al.
+2021a, b).
+UHD (Zheng et al.
+2021a) proposes a multi-guided bilateral upsampling model for UHD image dehazing.
+The core idea of it is to learn the local affine coefficients of the bilateral grid from the low-resolution image and apply them to the full-resolution image.
+UHDFour (Li et al.
+2023a) attempts to perform UHD low-light image enhancement in the Fourier domain and achieves good results on the real-world dataset.
+These methods are mainly based on CNN and overlook the modeling of global features.
+LLFormer (Wang et al.
+2023) and UHDformer (Wang et al.
+2024) both attempt to use transformers to capture the long-range dependency of features.
+The excellent long-range dependency capture capability of the transformer undoubtedly brings excellent results.
+However, LLFormer is also unable to perform full-resolution inference of UHD images on consumer-grade GPUs.
+It requires the image to be divided into multiple patches without overlapping for inference, and then the results are stitched together.
+This process undoubtedly degrades the quality of image restoration and may introduce boundary artifacts.
+While UHDFormer chooses to model the features globally in the 8$\times$ downsampled low-resolution space, this undoubtedly leads to the loss of too much high-frequency information, which is crucial for restoring the high-quality image.
+Furthermore, the inference time overhead introduced by the transformer architecture has been a persistent concern.
+In this paper, we propose a simple but effective UHD image restoration method, called MixNet.
 Specifically, we design the Feature Mixing Block which mainly contains Global Feature Modulation Layer (GFML), Local Feature Modulation Layer (LFML), and Feed-forward Layer (FFL).
 GFML is the key component of MixNet, enabling efficient modeling of long-range dependency.
-The use of self-attention in transformers endows them with powerful long-range dependency modeling capabilities Vaswani et al.
-(2017).
+The use of self-attention in transformers endows them with powerful long-range dependency modeling capabilities (Vaswani et al.
+2017).
 However, this approach comes with significant computational overhead.
-Inspired by the vanilla MLP-Mixer Tolstikhin et al.
-(2021), but different from it, GFML encodes the feature maps of images from the perspectives of the width $W$, height $H$, and channel $C$.
+Inspired by the vanilla MLP-Mixer (Tolstikhin et al.
+2021), but different from it, GFML encodes the feature maps of images from the perspectives of the width $W$, height $H$, and channel $C$.
 Through simple dimensional transformation operations, the information encoded in each feature map is associated and fused.
 It is worth noting that dimension transformation operations do not introduce additional parameters, which means that by permuting the feature maps from different views, GFML achieves efficient modeling of long-range dependency.
-LFML is designed to capture the local features of images.
+LFML is designed to capture the local features of the images.
 LFML adaptively calculates a set of parameters for the redistribution of weights in the channel dimension, emphasizing the importance of specific channels.
@@
 This way, our MixNet achieves a better trade-off between model complexity and performance.
-Extensive experiments on both synthetic and real-world datasets demonstrate that MixNet surpasses existing SOTA methods in terms of restoration quality and generalization ability.
+Extensive experiments on both synthetic and real-world datasets demonstrate that MixNet surpasses existing SOTA methods in terms of restoration quality and generalization ability (See Fig.
+1).
 Our contributions can be summarized below:
-  • We propose a simple but effective UHD-LLIE method called MixNet.
-MixNet can perform full-resolution inference of UHD images on consumer-grade GPUs.
-Compared to patch-based strategies, full-resolution inference can avoid the occurrence of boundary artifacts.
+  • We propose a simple but effective UHD image restoration method called MixNet.
+MixNet can perform full-resolution inference of UHD images on consumer-grade GPUs and achieve high-quality UHD image restoration.
   • We develop a novel efficient feature modulation mechanism to capture long-range dependency, which does not need to project the input image from 3D to 2D; it maintains the spatial properties between pixels by modeling directly on 3D space.
   • We evaluate the proposed method quantitatively and qualitatively on synthetic and real-world datasets, and the results show that our MixNet achieves a favorable trade-off between accuracy and model complexity.
-Furthermore, MixNet shows excellent scalability and flexibility, performing well on other image restoration tasks (e.g., dehazing).
-2 Related Work
-2.1 Low-Light Image Enhancement
-Learning-based methods have been successfully applied to the LLIE task Fu et al.
-(2022); Guo et al.
-(2020); Zhou et al.
-(2022); Wu et al.
-(2022); Xu et al.
-(2022).
-RetinexNet Wei et al.
-(2018) relies on the Retinex theory and designs a Decom-Net for decomposition and an Enhance-Net for illumination adjustment.
-Transformers have gained significant traction in computer vision since ViT Dosovitskiy et al.
-(2020) employs transformers for visual recognition tasks.
-Restormer Zamir et al.
-(2022) efficiently achieves image restoration by calculating self-attention at the channel dimension, thereby avoiding significant computational overhead.
-It demonstrates excellent performance across multiple tasks.
-Retinexformer Cai et al.
-(2023) introduces Retinex theory Land (1977) to the transformer and designs an illumination-guided transformer that utilizes illumination representations to direct the modeling of non-local interactions of regions with different lighting conditions.
-Diffusion models show realistic and detailed image generation through a sequence of denoising refinements.
-PyDiff Zhou et al.
-(2023) introduces diffusion models to LLIE to recover realistic details and performs sampling in a pyramid resolution style to get a fast generation.
-Although the methods above have achieved promising results in the LLIE task, they are not directly applicable to UHD images because the additional pixels of UHD images result in high computational overhead.
+Related Work
+UHD Image Restoration
+UHD image restoration has been an emerging topic in recent years (Liang, Zeng, and Zhang 2021; Zamfir, Conde, and Timofte 2023; Guan et al.
+2022).
+In  (Zheng et al.
+2021a, b), the authors introduced bilateral learning to achieve dehazing of the UHD image and HDR reconstruction of the UHD image.
+The core idea is to learn the local affine coefficients of the bilateral grid from the low-resolution image and apply them to the full-resolution image.
+UHD-SFNet (Wei, Zheng, and Jia 2022) and FourUHD (Li et al.
+2023a) explore underwater UHD image enhancement and UHD low-light image enhancement in the Fourier domain.
+They are all inspired by the discovery of most luminance information concentrates on amplitudes.
+Due to the unavoidable need for downsampling in existing UHD image restoration methods, NSEN (Yu et al.
+2023) proposes a spatial-variant and invertible nonuniform downsampler that adaptively adjusts the sampling rate according to the richness of details.
+LLFormer (Wang et al.
+2023) is the first transformer-based method to complete UHD-LLIE task, but like most previous methods, it is unable to perform full-resolution inference on consumer-grade GPUs.
+To avoid this issue, UHDformer (Wang et al.
+2024) designs two learning spaces, by building feature transformation from a high-resolution space to a low-resolution one, to compute self-attention in the low-resolution space, which significantly reduces the computational overhead.
+Despite the promising results achieved by the aforementioned approaches, there is still room for a favorable trade-off between reconstruction performance and model efficiency, prompting us to develop a more efficient global modeling method.
 Figure 2: The overall architecture of the proposed MixiNet.
-MixNet first transforms the input low-light image into the feature space using a DownSampler, performs feature extraction using a series of Feature Mixing Blocks (FMBs), and then reconstructs these extracted features using an UpSampler.
+MixNet first transforms the input degraded image into the feature space using a DownSampler, performs feature extraction using a series of Feature Mixing Blocks (FMBs), and then reconstructs these extracted features using an UpSampler.
 The FMB mainly contains a Global Feature Modulation Layer (GFML), a Local Feature Modulation Layer (LFML), and a Feed-forward Layer (FFL).
-2.2 UHD Image Restoration
-UHD image restoration has been an emerging topic in recent years Liang et al.
-(2021); Zamfir et al.
-(2023); Guan et al.
-(2022).
-In  Zheng et al.
-(2021a, b), the authors introduced bilateral learning to achieve dehazing of the UHD image and HDR reconstruction of the UHD image.
-The core idea is to learn the local affine coefficients of the bilateral grid from the low-resolution image and apply them to the full-resolution image.
-UHD-SFNet Wei et al.
-(2022) and FourUHD Li et al.
-(2023) explore underwater UHD image enhancement and UHD low-light image enhancement in the Fourier domain.
-They are all inspired by the discovery of most luminance information concentrates on amplitudes.
-Due to the unavoidable need for downsampling in existing UHD image restoration methods, NSEN Yu et al.
-(2023) proposes a spatial-variant and invertible nonuniform downsampler that adaptively adjusts the sampling rate according to the richness of details.
-LLFormer Wang et al.
-(2023) is the first transformer-based method to complete UHD-LLIE task, but like most previous methods, it is unable to perform full-resolution inference on consumer-grade GPUs.
-Despite the promising results achieved by the aforementioned approaches, there is still room for a favorable trade-off between reconstruction performance and model efficiency.
-3 Methodology
-MixNet aims to recover ideal UHD normal-light images from low-light images which often exhibit significantly reduced visibility, low contrast, and high levels of noise.
+Global Modeling
+In recent years, global modeling techniques have been receiving increasing attention in the field of computer vision, especially after the great success of Vision Transformer (Dosovitskiy et al.
+2020) in the visual domain.
+Transformer learns long-range dependencies between image patch sequences for global-aware modeling, and the interaction of these non-local features can significantly improve network performance.
+Various image restoration algorithms based on transformer have been proposed and achieve superior performance in restoration tasks such as image deraining (Chen et al.
+2023; Li et al.
+2023b), image dehazing (Song et al.
+2023; Qiu et al.
+2023), low-light image enhancement (Zhao et al.
+2021; Wu et al.
+2022; Xu et al.
+2022; Cai et al.
+2023) and image deblurring (Tsai et al.
+2022; Kong et al.
+2023).
+However, transformer-based methods are still difficult to apply directly to UHD image restoration tasks due to their huge computation cost.
+Owing to the low time complexity and hardware-friendly nature of MLP, some recent studies (Tolstikhin et al.
+2021) have explored the use of MLP for global modeling.
+However, MLP-based methods inevitably lose the spatial information between pixels during global modeling, which is unfavorable for pixel-level image restoration tasks.
+In contrast, our approach encodes the feature maps of images from the perspectives of width $W$, height $H$, and channel $C$, preserving spatial information while achieving efficient global modeling.
+Methodology
+MixNet aims to recover clean and realistic UHD images from degraded images which often exhibit significantly reduced visibility, low contrast, and high levels of noise.
 We provide the overall pipeline of our method and further details on the critical components of our approach below.
-3.1 Overview
+Overview
 An overview of MixNet is shown in Fig.
-2.
-For a given low-light image ${x}\in\mathbb{R}^{H\times W\times 3}$, we first map it to the feature space through a DownSampler to obtain the low-level features ${F}_{0}\in\mathbb{R}^{H\times W\times C}$, where $H$, $W$, and $C$ represent height, width, and channel, respectively.
+2 (a).
+For a given degraded image ${x}\in\mathbb{R}^{H\times W\times 3}$, we first map it to the feature space through a DownSampler to obtain the low-level features ${F}_{0}\in\mathbb{R}^{H\times W\times C}$, where $H$, $W$, and $C$ represent height, width, and channel, respectively.
 Then, the multiple stacked Feature Mixing Blocks (FMBs) are used to generate finer deep features ${F}_{i}$ from ${F}_{0}$ for normal-light image reconstruction, where a FMB has a Global Feature Modulation Layer (GFML), a Local Feature Modulation Layer (LFML) and a Feed-forward Layer (FFL).
-Lastly, the sum of the final features ${F}_{f}$ and ${F}_{0}$ is fed to a UpSampler to obtain the restored image ${\hat{y}}$.
-Figure 3: The overall architecture of the proposed (a) Global Feature Modulation Layer, (b) Local Feature Modulation Layer, and (c) Feed-forward Layer.
+Lastly, the sum of the final features ${F}_{f}$ and ${F}_{0}$ is fed to an UpSampler to obtain the restored image ${\hat{y}}$.
 Figure 3: Visualization of dimensional transformation operations. MixNet can capture long-range dependency of features with few parameters by employing simple dimension transformation operations.
-3.2 Global Feature Modulation Layer
-Recent studies suggest that the notable performance of transformers across diverse tasks stems from their implementation of the key multi-head self-attention (MHSA) mechanism Vaswani et al.
-(2017).
+Global Feature Modulation Layer
+Recent studies suggest that the notable performance of transformers across diverse tasks stems from their implementation of the key multi-head self-attention (MHSA) mechanism (Vaswani et al.
+2017).
 MHSA enables the model with the capability of long-range feature interaction and dynamic spatial weighting, both of which contribute to get promising results.
@@
 The architecture of the GFML is shown in Fig.
-3 (a) and the more details of dimension transformation operations are shown in Fig.
-4.
+2 (b) and the more details of dimension transformation operations are shown in Fig.
+3.
 Specifically, we first adjust the resolution of the normalized input features, and then perform some dimension transformation operations on them.
@@
 It is worth noting that before executing the last dimension transformation operation, we replace the GELU function with the Sigmoid function.
+Figure 4: Visual quality comparisons with state-of-the-art methods on UHD-LOL4K dataset. The last row shows the color histogram of the image. Our method has the closest color to GT image.
 Afterward, We use interpolation to adjust the feature $\hat{F}_{t}$ to its original resolution to estimate the attention map and adaptively modulate the input $F_{in}$ according to the estimated attention via element-wise product.
@@
 where $F_{g}$ are the final output features and $\odot$ represents element-wise product.
-3.3 Local Feature Modulation Layer
+Local Feature Modulation Layer
 LFML aims to capture the local feature of images and the architecture of it is shown in Fig.
-3 (b).
-For a set of feature maps, their impact on the final results is different Zhang et al.
-(2018b), and we hope the model can focus on the important feature maps.
+2 (c).
+For a set of feature maps, their impact on the final results is different (Zhang et al.
+2018), and we hope the model can focus on the important feature maps.
 We first feed the normalized input features into global average pooling to shift the focus of the model from space dimension to channel dimension, and then feed them into a series of convolutions to get deep features.
@@
 $\operatorname{Sigmoid}(\cdot)$ represents Sigmoid function, $\operatorname{Conv}(\cdot)$ is a $1\times 1$ convolution and $\operatorname{GAP}(\cdot)$ denotes global average pooling.
-3.4 Feed-forward Layer
+Feed-forward Layer
+Through GFML and LFML, we obtain global features and local features respectively.
+Undoubtedly, there will be overlapping parts in the feature space between these two, leading to redundant information.
 To transform features into a compact representation, we introduce FFL into our model.
 As shown in Fig.
-3 (c), FFL comprises a $3\times 3$ convolution, a $1\times 1$ convolution, and a GELU function.
+2 (d), FFL comprises a $3\times 3$ convolution, a $1\times 1$ convolution, and a GELU function.
 Within this, the first $3\times 3$ convolution encodes the spatially local contexts and doubles the number of channels of the input features for mixing channels; the later $1\times 1$ convolution reduces the channels back to the original input dimension.
-In this way, the model learns a more compact representation of features and ignores some unimportant information Tan and Le (2021).
+In this way, the model learns a more compact representation of features and ignores some unimportant information (Tan and Le 2021).
 The FFL can be expressed as:
@@
 where $F_{in}$ and $F_{out}$ are the input features and output features, respectively.
-3.5 Feature Mixing Block
+Feature Mixing Block
 Here, we describe our FMB pipeline in general, which can be written as:
@@
 [;] is concatenation operation, and $\operatorname{Conv}$ respresents the $1\times 1$ convolution.
-3.6 Loss Functions
+Loss Functions
 To optimize the weights and biases of the network, we utilize the L1 loss in the RGB color space as the basic reconstruction loss:
@@
 where $y$ and $\hat{y}$ denote the ground truth and restored image, respectively.
-Figure 5: Visual quality comparisons. The image can be zoomed in for better visualization. The last row shows the color histogram of the image. Our method has the closest color to GT image.
-Table 1: Comparison of quantitative results on UHD-LOL4K and UHD-LL datasets. Best and second best values are indicated with bold text and underlined text respectively.
-Type Methods UHD-LOL4K UHD-LL Average Parameter
-PSNR SSIM LPIPS PSNR SSIM LPIPS PSNR SSIM LPIPS
-non-UHD Zero-DCE Guo et al. (2020) 17.19 0.850 0.193 17.08 0.663 0.513 17.14 0.757 0.353 79.416K
-Zero-DCE++ Li et al. (2021) 15.58 0.835 0.222 16.41 0.630 0.530 16.00 0,733 0.376 10.561K
-RUAS Liu et al. (2021) 14.68 0.758 0.274 13.56 0.749 0.460 14.12 0.754 0.367 3.438K
-Uformer Wang et al. (2022) 29.99 0.980 0.034 19.28 0.849 0.356 24.64 0.915 0.195 20.628M
-Restormer Zamir et al. (2022) 36.91 0.988 0.023 22.25 0.871 0.289 29.58 0.930 0.156 26.112M
-UHD LLFormer Wang et al. (2023) 37.33 0.989 0.020 22.79 0.853 0.264 30.06 0.921 0.142 13.152M
-UHDFour Li et al. (2023) 36.12 0.990 0.021 26.22 0.900 0.239 31.17 0.945 0.130 17.537M
-MixNet 39.22 0.992 0.006 27.54 0.862 0.134 33.38 0.927 0.070 7.770M
-4 Experiment
-4.1 Implementation Details and Datasets
-4.1.1 Implementation Details
+Figure 5: Visual quality comparisons with state-of-the-art methods on UHD-UW dataset. Please zoom in for details.
+Methods Type Venue PSNR $\uparrow$ SSIM $\uparrow$ Param $\downarrow$
+Z_DCE++ non-UHD TPAMI’21 15.58 0.934 10.56K
+RUAS CVPR’21 14.68 0.757 3.44K
+Uformer CVPR’22 29.98 0.980 20.63M
+Restormer CVPR’22 36.90 0.988 26.11M
+NSEN UHD MM’23 29.49 0.980 2.67M
+UHDFour ICLR’23 36.12 0.990 17.54M
+LLFormer AAAI’23 37.33 0.988 24.52M
+UHDformer AAAI’24 36.28 0.989 339.3K
+MixNet - 39.22 0.992 7.77M
+Table 1: Comparison of quantitative results on UHD-LOL4K dataset. Best and second best values are indicated with bold text and underlined text respectively.
+Experiments
+Implementation Details and Datasets
+Implementation Details.
 We conduct experiments in PyTorch on six NVIDIA GeForce RTX 3090 GPUs.
-To optimize the network, we employ the Adam optimizer with a learning rate $2\times 10^{-4}$.
-We randomly crop the full-resolution 4K image to a resolution of $512\times 512$ as the input and perform 300k iterations of training with a batch size of 24.
+To optimize the network, we employ the Adam optimizer with a initial learning rate $2\times 10^{-4}$ and a cosine annealing strategy is used for the decay of the learning rate.
+We randomly crop the full-resolution 4K image to a resolution of $512\times 512$ as the input and the batch size is set ot 24.
+Except for the UHD underwater image enhancement task, where the dataset is relatively small, we only perform 100K iterations.
+For all other tasks, we perform 300K iterations.
 To augment the training data, random horizontal and vertical flips are applied to the input images.
 The number of FMB and feature channels is set to 8 and 48, respectively.
-DownSampler and UpSampler are both composed of a sub-pixel convolutional layer Shi et al.
-(2016) and a $3\times 3$ convolutional layer.
-4.1.2 Datasets
-UHD-LOL.
-UHD-LOL Wang et al.
-(2023) is a large-scale benchmark dataset containing UHD images in low-light conditions and includes two subsets, UHD-LOL4K and UHD-LOL8K, containing 4K and 8K-resolution images, respectively.
-In this paper, we use the UHD-LOL4K to verify the effectiveness of MixNet.
-UHD-LOL4K contains 8,099 image pairs, 5,999 for training and 2, 100 for testing.
-UHD-LL.
-UHD-LL Li et al.
-(2023) is a real-world low-noise/normal-clear paired image dataset that contains 2,150 pairs of 4K UHD data saved in 8bit sRGB format.
-It has been split into two parts: 2,000 pairs for training and 115 pairs for testing.
-LOL.
-In addition to the aforementioned UHD datasets, we also evaluate our method on the LOL dataset Wei et al.
-(2018) which is the commonly used dataset in the LLIE task.
-LOL has 500 image pairs and the training and testing sets are split in proportion to 485:15.
-Figure 6: Visual quality comparisons. The image can be zoomed in for improved visualization. Our method can effectively remove the noise and reconstruct well-exposed image details.
-4.2 Comparisons with the State-of-the-art Methods
-4.2.1 Quantitative Results
-We employ three well-known image quality assessment metrics, PSNR, SSIM Wang et al.
-(2004) and LPIPS (Alex version) Zhang et al.
-(2018a) to quantify the performance of different methods.
+Datasets.
+To verify the effectiveness of MixNet, we evaluate it on the UHD-LOL dataset (Wang et al.
+2023), UHD-Blur dataset (Wang et al.
+2024), UHD-UW dataset (Wei, Zheng, and Jia 2022) and UHDM dataset (Yu et al.
+2022a).
+UHD-LOL consists of two subsets, namely UHD-LOL4K and UHD-LOL8K, which contain UHD images with resolutions of 4K and 8K in low-light conditions, respectively.
+In this study, we focus on the UHD-LOL4K subset to validate the effectiveness of WowNet.
+The UHD-LOL4K subset comprises a total of 8,099 image pairs, with 5,999 pairs allocated for training purposes and 2,100 pairs designated for testing.
+UHD-Blur dataset is re-collected from the datasets of Deng et al.
+(Deng et al.
+2021) by Wang et al.
+(Wang et al.
+2024) and it has 1,964 blur-clear image pairs for training and 300 pairs for testing.
+UHD-UW dataset is our organization of the dataset proposed in UHD-SFNet (Wei, Zheng, and Jia 2022), where we remove the images with resolutions lower than $512\times 512$, and sequentially select 80 images as our training set, with the remaining 13 images as our test set.
+UHDM dataset is a real-world 4K resolution demoir´eing dataset, consisting of 5,000 image pairs in total, with 4500 images used for training and 500 images used for testing.
+Methods Type Venue PSNR $\uparrow$ SSIM $\uparrow$ Param $\downarrow$
+Shallow-uwnet non-UHD AAAI’21 23.36 0.798 0.22M
+USUIR AAAI’22 24.13 0.834 3.36M
+PUIE-Net ECCV’22 25.77 0.859 1.40M
+URank AAAI’23 26.39 0.911 3.15M
+FA+Net BMVC’23 27.60 0.916 9K
+UHD UHD ICCV’21 26.89 0.907 34.5M
+UHD-SFNet ACCV’22 26.31 0.914 37.31M
+Ours - 28.09 0.932 7.77M
+Table 2: Comparison of quantitative results on UHD-UW dataset. Best and second best values are indicated with bold text and underlined text respectively.
+Comparisons with the State-of-the-art Methods
+Evaluation.
+Two well-known metrics, Peak Signal-to-Noise Ratio (PSNR) and Structural Similarity (SSIM) (Wang et al.
+2004), are employed for quantitative comparisons.
+Higher values of these metrics indicate superior performance of the methods.
 PSNR and SSIM are calculated along the RGB channels.
-For UHD datasets, we compare our method with Zero-DCE Guo et al.
-(2020), Zero-DCE++ Li et al.
-(2021), RUAS Liu et al.
-(2021), Uformer Wang et al.
-(2022), Restormer Zamir et al.
-(2022), LLFormer Wang et al.
-(2023) and UHDFour Li et al.
-(2023).
-It is worth noting that most methods cannot directly perform full-resolution inference on UHD images.
-Therefore, for these methods, we adopt a strategy of dividing the images into multiple patches without overlapping for interference, and then the results are stitched together.
-For the LOL dataset, in addition to the mentioned methods, we also compare our MixNet with methods RetinexNet Wei et al.
-(2018), KinD Zhang et al.
-(2019), DSLR Lim and Kim (2020), RUAS Liu et al.
-(2021), ELGAN Jiang et al.
-(2021) and Retinexformer Cai et al.
-(2023).
-In Tab.
-1, we report the quantitative comparison results among different methods on the UHD dataset.
-When compared with the previous SOTA method FourUHD, our MixNet gains by an average of 2.21dB on the UHD-LOL4K and UHD-LL datasets.
-Compared to other methods specifically designed for UHD images, our MixNet achieves impressive results with few parameters.
-Our method also performs well on the LOL dataset and the results are reported in Tab.
-2.
-It is worth noting that the resolution of LOL images is relatively low.
-Therefore, we replace the sub-pixel convolutional layer in DownSampler and UpSampler with the Residual Block of comparable parameters.
-All these results clearly suggest the outstanding effectiveness and efficiency advantage of our MixNet.
-4.2.2 Qualitative Results
-In addition to the quantitative evaluations, we also provide qualitative comparisons of the proposed method.
-Figs.
-5 and  6 present the visual comparisons.
-MixNet can effectively enhance the poor visibility and low contrast or low-light regions, reliably remove the noise without introducing spots and artifacts, and robustly preserve the color.
-Table 2: Comparison of quantitative results on LOL dataset. Best and second best values are indicated with bold text and underlined text respectively.
-Type Methods PSNR SSIM
-non-UHD RetinexNet Wei et al. (2018) 16.77 0.425
-KinD Zhang et al. (2019) 17.65 0.772
-DSLR Lim and Kim (2020) 14.98 0.596
-Z_DCE Guo et al. (2020) 14.86 0.562
-Z_DCE++ Li et al. (2021) 14.75 0.518
-RUAS Liu et al. (2021) 16.40 0.503
-ELGAN Jiang et al. (2021) 17.48 0.652
-Uformer Wang et al. (2022) 18.55 0.721
-Restormer Zamir et al. (2022) 22.37 0.816
-Retinexformer Cai et al. (2023) 25.16 0.845
-UHD UHDFour Li et al. (2023) 23.09 0.871
-LLFormer Wang et al. (2023) 23.65 0.816
-MixNet 24.21 0.860
-4.3 Ablation Study
+We also report the trainable parameters (Param) and running time (RT) which you can find in Tab.
+7.
+For methods that cannot perform full-resolution inference, we adopt the previous patch-based or downsample-based approaches to handle them.
+Figure 6: Visual quality comparisons with state-of-the-art methods on UHD-Blur dataset. Please zoom in for details.
+Methods Type Venue PSNR $\uparrow$ SSIM $\uparrow$ Param $\downarrow$
+MIMO-Unet++ non-UHD ICCV’21 25.03 0.752 16.1M
+Restormer CVPR’22 25.21 0.752 26.1M
+Uformer CVPR’22 25.27 0.752 20.6M
+Stripformer ECCV’22 25.05 0.750 19.7M
+FFTformer CVPR’23 25.41 0.757 16.6M
+UHDformer UHD AAAI’24 28.82 0.844 0.34M
+Ours - 29.43 0.855 5.22M
+Table 3: Comparison of quantitative results on UHD-Blur dataset. Best and second best values are indicated with bold text and underlined text respectively.
+Low-light UHD Image Enhancement Results.
+For low-light UHD image enhancement task, we compare our MixNet with approaches Z_DCE++ (Li, Guo, and Loy 2021), RUAS (Liu et al.
+2021), ELGAN (Jiang et al.
+2021), Uformer (Wang et al.
+2022), Restormer (Zamir et al.
+2022), NSEN (Yu et al.
+2023), UHDFour (Li et al.
+2023a), LLFormer (Wang et al.
+2023) and UHDformer (Wang et al.
+2024).
+As shown in Tab.
+1, our method outperforms the current state-of-the-art method, LLFormer, by 1.89dB in PSNR, which is undoubtedly a significant improvement.
+And you can find the visualization results in Fig.
+4.
+From the color histograms, it can be seen that the images restored using our method have the color distribution closest to the ground truth.
+Underwater UHD Image Enhancement Results.
+We evaluate underwater UHD image enhancement results on UHD-UW dataset and compare our method with several methods (Naik, Swarnakar, and Mittal 2021; Fu et al.
+2022a, b; Guo et al.
+2023; Jiang et al.
+2023; Zheng et al.
+2021a; Wei, Zheng, and Jia 2022).
+The results are shown in Tab.
+5, and compared with recent state-of-the-art FA+Net (Jiang et al.
+2023), MixNet achieves a 0.49dB performance improvement in PSNR.
+You can see the comparison of the visualization results in Fig.
+5, which shows that our method achieves better color correction.
+UHD Image Deblurring Results.
+In the UHD image deblurring task, we evaluate our proposed MixNet against existing deblurring approaches, including MIMO-Unet++ (Cho et al.
+2021), Restormer (Zamir et al.
+2022), Uformer (Wang et al.
+2022), Stripformer (Tsai et al.
+2022), FFTformer (Kong et al.
+2023) and UHDformer (Wang et al.
+2024).
+As shown in Tab.
+3, our method achieved a 0.61 dB improvement in PSNR compared to the current state-of-the-art method, UHDformer.
+Additionally, you can find the visual comparison results in Fig.
+6.
+UHD Image Demoiréing Results.
+To validate the effectiveness of our method on the task of moiré pattern removal, we compare it with many methods, including MDDM (Cheng, Fu, and Yang 2019), MopNet (He et al.
+2019), WDNet (Liu et al.
+2020), MBCNN (Zheng et al.
+2020), FHDe2Net (He et al.
+2020) and ESDNet (Yu et al.
+2022b).
+As shown in Tab.
+4, our method achieves a 0.23 dB improvement in PSNR compared to the ESDNet, which is also specifically designed for UHD image demoiréing.
+You can find the comparison of visualization results in Fig.
+7.
+Methods Type Venue PSNR $\uparrow$ SSIM $\uparrow$ Param $\downarrow$
+MDDM non-UHD ICCVW’19 20.09 0.744 7.63M
+MopNet ICCV’19 19.48 0.757 58.5M
+WDNet ECCV’20 20.36 0.649 3.36M
+MBCNN CVPR’20 21.41 0.793 14.19M
+FHDe2Net ECCV’20 20.33 0.749 13.5M
+ESDNet UHD ECCV’22 22.11 0.795 5.93M
+Ours - 22.34 0.799 7.77M
+Table 4: Comparison of quantitative results on UHDM dataset. Best and second best values are indicated with bold text and underlined text respectively.
+Figure 7: Visual quality comparisons with state-of-the-art methods on UHDM dataset. Please zoom in for details.
+Ablation Study
 We further conduct extensive ablation studies to better understand and evaluate each component in the proposed MixNet.
 For fair comparisons with the designed baselines, we implement all experiments with the same setting on UHD-LOL4K dataset.
-We gradually replace the proposed module with the Residual Block (ResBlock) He et al.
-(2016) of comparable parameters, and the experimental results in Tab.
-4.
-Based on all the results in the Tab.
-4, all the key designs contribute to the best performance of the full model.
-Without the GFML (#a, #c, #d, #f), the model shows the most severe performance degradation.
-This aligns with our hypothesis, as the excellent long-range dependency modeling capability brought by GMFL is one of the key factors contributing to the outstanding performance of MixNet.
-And the result suggests the importance of long-range dependency modeling among features.
-From the results of #g and #e, we find that the impact of FFL on performance is higher than that of GFML.
-We believe that this might be the ability of the FFL to learn compact feature representations, which the ResBlock lacks.
-And any CNN is capable of learning local features.
-From the results of #c and #d, we observe that when using the block we proposed alone, LFML has a higher impact on the results than FFL.
-This may be because when the feature representation is not effective enough (not from GFML and LFML), learning its compact representation has little effect.
-We also provide visualizations of the feature maps corresponding to #e, #f, #g and #h.
-As shown in Fig.
-7, it is evident that our method reconstructs feature maps with sharper textures.
-Table 3: Comparison of quantitative results on O-Haze dataset. Best and second best values are indicated with bold text and underlined text respectively.
-Task Methods PSNR SSIM
-non-UHD DCP He et al. (2010) 14.54 0.333
-BCCR Meng et al. (2013) 15.20 0.361
-CAP Zhu et al. (2015) 15.35 0.342
-NLD Berman et al. (2016) 14.76 0.396
-DehazeNet Cai et al. (2016) 12.39 0.359
-AodNet Li et al. (2017) 15.87 0.581
-GridNet Liu et al. (2019) 17.44 0.603
-PMSNet Chen et al. (2019) 16.99 0.597
-MSBDN Dong et al. (2020) 17.46 0.635
-UHD MGBL Zheng et al. (2021a) 18.72 0.614
-NSEN Yu et al. (2023) 19.23 0.672
-MixNet 21.46 0.650
-4.4 Extension Experiments
-Although we primarily focus on the UHD-LLIE task in this paper, it is important to note that we do not introduce specific prior knowledge related to low-light conditions.
-Therefore, our method can be considered as a general approach that performs well not only in UHD-LLIE task but also in various UHD image restoration task, such dehazing.
-4.4.1 Datasets
-We conduct dehazing experiments on the O-HAZE dataset which has 45 paired hazy and haze-free images.
-Following the prior work Yu et al.
-(2023), we randomly choose 35 images for training and the others for testing.
-4.4.2 Results
-We compare the proposed MixNet with several image dehazing approaches Berman et al.
-(2016); He et al.
-(2010); Meng et al.
-(2013); Zhu et al.
-(2015); Li et al.
-(2017); Dong et al.
-(2020); Liu et al.
-(2019); Cai et al.
-(2016); Chen et al.
-(2019); Zheng et al.
-(2021a); Yu et al.
-(2023).
-Results are reported in Tab.
-3.
-MixNet achieves a substantial gain of over 2 dB compared to NSEN in PSNR on the UHD image dehazing task.
-Table 4: Ablation study of proposed blocks. In order to ensure consistency in parameter magnitude, we gradually replace the proposed module with Residual Block of comparable parameters.
-# GFML LFML FFL PSNR SSIM Parameter
+# GFML LFML FFL PSNR $\uparrow$ SSIM $\uparrow$ Param $\downarrow$
 (a) 30.73 0.968 7.178M
@@
 (h) 39.22 0.992 7.770M
-Figure 7: Visualizations of the feature maps before the UpSampler. The image used here is the low-light image shown in Fig. 6. Our method reconstructs feature maps with sharper textures.
-5 Conclusion
-In this paper, we proposed a simple yet efficient framework called MixNet to solve the efficient UHD-LLIE problem.
+Table 5: Ablation study of proposed blocks. In order to ensure consistency in parameter magnitude, we gradually replace the proposed module with ResBlock.
+Stage1 Stage2 Stage3
+PSNR $\uparrow$ 37.94 38.57 39.22
+SSIM $\uparrow$ 0.989 0.990 0.992
+Table 6: Ablation study of the number of permute stages. We gradually remove some stages to verify their necessity.
+Effectiveness of Proposed Modules.
+To demonstrate the gains of proposed modules, we gradually replace the proposed modules with the Residual Block (ResBlock) (He et al.
+2016) of comparable parameters, and the experimental results in Tab.
+5.
+Based on all the results in the Tab.
+5, all the key designs contribute to the best performance of the full model.
+Without the GFML (#a, #c, #d, #f), the model shows the most severe performance degradation.
+This aligns with our hypothesis, as the excellent long-range dependency modeling capability brought by GMFL is one of the key factors contributing to the outstanding performance of MixNet.
+And the result suggests the importance of long-range dependency modeling among features.
+From the results of #g and #e, we find that the impact of FFL on performance is higher than that of GFML.
+We believe that this might be the ability of the FFL to learn compact feature representations, which the ResBlock lacks.
+And any CNN is capable of learning local features.
+From the results of #b, #c and #d, we observe that when using the block we proposed alone, GFML and LFML have a higher impact on the results than FFL.
+This may be because when the feature representation is not effective enough (not from GFML and LFML), learning its compact representation has little effect.
+Effectiveness of the Number of Permute Stages.
+GFML is the core part of our method.
+As can be clearly seen in Fig.
+3, its main process can be divided into three steps.
+Here we verify the necessity of these three stages.
+In Tab 6, we can see that the network’s performance drops sharply when some stages are removed.
+We believe this is because when part of the permute operation is removed, the global modeling capability of GFML is affected, and the remaining permute operation is not enough to complete the long-range dependency modeling of the features.
+UHD LLFormer UHDFour UHDformer Ours
+Param $\downarrow$ 34.5M 24.52M 17.54M 0.34M 7.70M
+RT $\downarrow$ 1.58s 7.12s 1.67s 5.85s 0.53s
+Table 7: Trainable parameters and running time of some UHD image restoration methods.
+We record the time cost required to perform inference on $3840\times 2160$ size images using a single NVIDIA GeForce RTX 3060 GPU.
+Figure 8: Examples of unsatisfactory restorations. Please zoom in for details.
+Conclusion
+In this paper, we propose a simple yet effective framework called MixNet to solve the efficient UHD image restoration problem.
 MixNet has a series of FMBs which mainly contain a GFML, LFML, and FFL.
-GFML explored long-range dependency modeling upon the simple dimensional transformation operation.
+GFML explores long-range dependency modeling upon the simple dimensional transformation operation.
 By permuting the feature maps, GFML associates features from different views.
@@
 Extensive experiments on both synthetic and real-world datasets showed that the proposed MixNet is more efficient than state-of-the-art methods while achieving competitive performance.
+Limitations.
+Although our proposed MixNet achieves satisfactory performance across multiple tasks, its efficacy remains limited when dealing with real-world datasets containing complex degradation patterns.
+For example, in the task of moiré pattern removal, as shown in Fig.
+8, the complex moire patterns have not been effectively removed, and the visual results still have a certain gap from the ground truth.